# Macroscopic Quantum States and Universal Correlations in a Disorder-Order Interface Propagating over a 1D Ground State

### v2

@@ Abstract
We compute $I_{\rho_{A_{t}}}(\boldsymbol{\bf S}^{x})$ numerically, for $A_{t}$ in the edge comoving frame. We find $\lim_{t\rightarrow\infty}t^{-2/3}I_{\rho_{A_{t}}}(\boldsymbol{\bf S}^{x})\neq 0$. Thus, subsystems of size comparable with the interfacial region are in macroscopic quantum states. Figure 3 reports some numerical data compared with a semiclassical approximation, detailed in Ref. [75]. Although the quantum contribution to the variance is small in comparison to the classical part, $I_{\rho_{A_{t}}}(\boldsymbol{\bf S}^{x})$ still witnesses quantum correlations over a distance of order $t^{1/3}$.
@@ Appendix A End Matter
$\begin{pmatrix}\mathcal{H}_{2j-1,2\ell-1}&\mathcal{H}_{2j-1,2\ell}\\ \mathcal{H}_{2j,2\ell-1}&\mathcal{H}_{2j,2\ell}\end{pmatrix}=\int_{-\pi}^{\pi}\frac{dp}{2\pi}\ h(e^{ip})e^{ip(j-\ell)}$
@@
$h(e^{ip})=\varepsilon(p)e^{-i\frac{\theta(p)}{2}\sigma^{z}}\sigma^{y}e^{i\frac{\theta(p)}{2}\sigma^{z}}\;,$
@@
$\begin{pmatrix}\Gamma_{2j-1,2\ell-1}&\Gamma_{2j-1,2\ell}\\ \Gamma_{2j,2\ell-1}&\Gamma_{2j,2\ell}\end{pmatrix}=\int_{-\pi}^{\pi}\frac{dp}{2\pi}\ \Gamma(e^{ip})e^{ip(j-\ell)}\,.$
@@
$\Gamma_{2j-2+i,2\ell-2+i^{\prime}}=\int_{-\pi}^{\pi}\frac{dp}{2\pi}\ [\Gamma_{\frac{j+\ell}{2}}]_{i,i^{\prime}}(e^{ip})e^{ip(j-\ell)}\;,$
@@
$\partial_{t}\rho_{x,t}^{(3)}(p)+v(p)\partial_{x}\rho^{(3)}_{x,t}(p)=\tfrac{v^{\prime\prime}(p)}{24}\partial_{x}^{3}\rho^{(3)}_{x,t}(p)\ ,$
where $v(p)=\frac{d\varepsilon(p)}{dp}$ is the velocity of the quasiparticle excitation with momentum $p$. The third-order correction is particularly important close to the edge of the lightcone, where it gives a leading contribution. In a bipartitioning protocol where one side of the system, e.g., the right-hand side, is initially prepared in a stationary state, the auxiliary field diffuses over the right hand side of the junction and therefore has a negligible effect along any ray with a strictly positive slope.

### v3

@@ Abstract
We compute $I_{\rho_{A_{t}}}(\boldsymbol{\bf S}^{x})$ numerically, for $A_{t}$ in the edge comoving frame with free fermion techniques [78]. We find $\lim_{t\rightarrow\infty}t^{-2/3}I_{\rho_{A_{t}}}(\boldsymbol{\bf S}^{x})\neq 0$. Thus, subsystems of size comparable with the interfacial region are in macroscopic quantum states. Figure 3 compares numerical data with a semiclassical approximation, detailed in Ref. [78]. Although the quantum contribution to the variance is small in comparison to the classical part, $I_{\rho_{A_{t}}}(\boldsymbol{\bf S}^{x})$ witnesses quantum correlations over a distance of order $t^{1/3}$.
@@ Appendix A End Matter
$\begin{pmatrix}\mathcal{H}_{2j-1,2\ell-1}&\mathcal{H}_{2j-1,2\ell}\\ \mathcal{H}_{2j,2\ell-1}&\mathcal{H}_{2j,2\ell}\end{pmatrix}=\int_{-\pi}^{\pi}\frac{dp}{2\pi}\ h(e^{ip})e^{ip(j-\ell)}$ (10)
@@
$h(e^{ip})=\varepsilon(p)e^{-i\frac{\theta(p)}{2}\sigma^{z}}\sigma^{y}e^{i\frac{\theta(p)}{2}\sigma^{z}}\;,$ (11)
@@
$\begin{pmatrix}\Gamma_{2j-1,2\ell-1}&\Gamma_{2j-1,2\ell}\\ \Gamma_{2j,2\ell-1}&\Gamma_{2j,2\ell}\end{pmatrix}=\int_{-\pi}^{\pi}\frac{dp}{2\pi}\ \Gamma(e^{ip})e^{ip(j-\ell)}\,.$ (12)
@@
$\Gamma_{2j-2+i,2\ell-2+i^{\prime}}=\int_{-\pi}^{\pi}\frac{dp}{2\pi}\ [\Gamma_{\frac{j+\ell}{2}}]_{i,i^{\prime}}(e^{ip})e^{ip(j-\ell)}\;,$ (13)
@@
$\partial_{t}\rho_{x,t}^{(3)}(p)+v(p)\partial_{x}\rho^{(3)}_{x,t}(p)=\tfrac{v^{\prime\prime}(p)}{24}\partial_{x}^{3}\rho^{(3)}_{x,t}(p)\ ,$ (14)
where $v(p)=\frac{d\varepsilon(p)}{dp}$ is the velocity of the quasiparticle excitation with momentum $p$, and the superscript $(3)$ is a reminder that $\rho$ is equal to $\rho^{(3)}$ only up to subleading contributions—in the limit of infinite time—coming from spatial derivatives higher than the third. The third-order correction is particularly important close to the edge of the lightcone, where it gives a leading contribution. In a bipartitioning protocol where one side of the system, e.g., the right-hand side, is initially prepared in a stationary state, the auxiliary field is instead characterized by rapid oscillations and an amplitude that approaches zero as $t^{-\frac{1}{2}}$ along any ray with a finite slope strictly within the lightcone, and as $t^{-2/3}$ at the edge of the lightcone [78].